# Grid-based Multi-Robot Coverage Path Planning

### v1

@@ 2 $M\leftarrow$ find MST of $H$ via Kruskal’s algorithm
3 $\pi\leftarrow$ compute the circumnavigating path of $M$ with the Full-STC routing rules via Depth-First-Search
4 return the coverage path $\pi$
@@ IV-B Theoretical Analysis
This subsection shows that ESTC achieves complete coverage in Theorem 3 and is the optimal circumnavigating coverage path in Theorem 4. We demonstrate that ESTC is bounded suboptimal for CPP (including non-circumnavigating paths) in Theorem 5. We further analyze the time complexity of ESTC in Theorem 6, highlighting its practical time efficiency.
@@
ESTC achieves complete coverage on a connected graph $G$.
Proof 3.
The proof is analogous to the argument used in the proof of Lemma 3.1 in [gabriely2002spiral] given that $G$ is connected.
@@
The cost of the circumnavigating path from a minimum spanning tree of the hypergraph is always smaller than that from any non-minimum spanning tree.
Proof 4.
For hypergraph $H=(V_{h},E_{h})$, consider a minimum spanning tree $M=(V_{h},F\cup\{\varepsilon_{m}\})$ of $H$ and a non-minimal spanning tree $N=(V_{h},F\cup\{\varepsilon_{n}\})$ of $H$. Denote $c(\pi_{M})$ and $c(\pi_{N})$ as the costs of the two circumnavigating paths resulting from $M$ and $N$, respectively. As illustrated in Fig. 4, we know that $E(\pi_{M})\setminus E(\pi_{N})=\{(u_{1},v_{1}),(u_{2},v_{2})\}$ (blue solid lines) and $E(\pi_{N})\setminus E(\pi_{M})=\{(u_{1},u_{2}),(v_{1},v_{2})\}$ (red solid lines) if $x_{\varepsilon_{m}}=x_{\varepsilon_{n}}=2$. Therefore, we can compute the cost difference as $\Delta c=c(\pi_{M})-c(\pi_{N})$ as follows:
$\Delta c$ $=[w_{(u_{1},v_{1})}+w_{(u_{2},v_{2})}-w_{(u_{1},u_{2})}-w_{(v_{1},v_{2})}]$
$-[w_{(p_{1},q_{1})}+w_{(p_{2},q_{2})}-w_{(p_{1},q_{2})}-w_{(p_{1},q_{2})}]$
Substituting Eqn. (2) into the above equation, we have $\Delta c=w_{\varepsilon_{m}}-w_{\varepsilon_{n}}<0$ as $w_{\varepsilon_{m}}<w_{\varepsilon_{n}}$ by definition. As a result, we have $c(\pi_{M})<c(\pi_{N})$ and the same derivation applies to all other combinations of $(x_{\varepsilon_{m}},x_{\varepsilon_{n}})\in\{1,2\}\times\{1,2\}$. By replacing an edge from $M$ to an edge from an arbitrary non-minimum spanning tree $T$ one at a time, we can sequentially prove $c(\pi_{M})<c(\pi_{T})$ using the above derivation and thus conclude the proof.
@@
ESTC is bounded suboptimal for CPP.
Proof 5.
@@
Considering the CPP problem for the MRS $(G,I,R)$, we analyze the worst-case cost of solution $\pi$ by ESTC compared with the optimal CPP solution $\pi^{*}$. Suppose that graph $G$ is contracted into hypergraph $H=(V_{h},E_{h})$ with $n_{i}$ incomplete hypervertices. To ensure that $\pi^{*}$ is a valid complete coverage path, we know that $V(\pi)=V$ and the first and last vertices of $\pi$ are both the root vertex. It follows that one can easily derive $|E(\pi^{*})|\geq|V|$, where the equality holds iff $n_{i}=0$. Then we can obtain the optimality ratio $\rho$ for ESTC:
$\rho=\frac{c(\pi)}{c(\pi^{*})}=\frac{\sum_{e\in E_{\pi}}}{\sum_{e\in E_{\pi}^{*}}}\leq\frac{|E_{\pi}|\cdot w_{\text{max}}}{|E_{\pi}^{*}|\cdot w_{\text{min}}}\leq\frac{|E_{\pi}|\cdot w_{\text{max}}}{|V|\cdot w_{\text{min}}}$ (3)
where $w_{\text{max}}$ and $w_{\text{min}}$ are the maximal and minimal weights among all hyperedges from $E_{h}$, respectively. If $n_{i}=0$, we have $|E_{\pi}|=|V|$ following the same derivation as above for $\pi^{*}$. If $n_{i}\neq 0$, ESTC adds at most two additional edges that duplicate with existing edges compared to the case of $n_{i}=0$, to traverse each incomplete hypervertex. As a result, we have $|E(\pi)|\leq|V|+2\cdot n_{i}$, with which we can further refine Eqn. (3) as follows:
$\rho\leq(1+2\cdot\frac{n_{i}}{|V|})\cdot\frac{w_{\text{max}}}{w_{\text{min}}}$ (4)
Clearly, the right-hand side of Eqn. (4) depends only on the input $G$, implying that ESTC is bounded suboptimal.
@@
The above bound of the optimality ratio is rather loose. Since $n_{i}\in[0,|V|]$, we know that the first multiplier $(1+2\cdot\frac{n_{i}}{|V|})\in[1,3]$. In practice, when $n_{i}$ is large, the optimal solution $\pi^{*}$ must adopt rerouting rules (e.g., Full-STC) to revisit edges through incomplete hypervertices for complete coverage. This puts $|E(\pi^{*})|$ close to $|E(\pi)|$ instead of $|V|$, which narrows the bound in Eqn. (3). Specifically, ESTC is optimal ($\rho=1$) if $G$ is unweighted ($\frac{w_{\text{max}}}{w_{\text{min}}}=1$) and all hypervertices are complete ($n_{i}=0$).
@@
The time complexity of ESTC (Alg. 1) for any grid graph $G=(V,E)$ is $O(|V|\log|V|)$.
Proof 6.
To construct hypergraph $H=(V_{h},E_{h})$ from $G=(V,E)$ as described in Sec. IV-A, we iterate $V$ to build $V_{h}$ and then iterate $V_{h}$ to build $E_{h}$. As $|V_{h}|$ is proportional to $|V|$, the time complexity of this construction is $O(|V|)$. To find the MST of $H$, ESTC uses Kruskal’s algorithm therefore having a time complexity of $O(|E|\log|E|)$. The procedure to generate the circumnavigating coverage path on the MST via Depth-First-Search entails a time complexity of $O(|V|+|E|)$. As $|E|$ is proportional to $|V|$ in the grid graph $G$, combining the above, the time complexity of ESTC is $O(|V|\log|V|)$.
IV-C Local Optimizations
We introduce two efficient local optimizations for ESTC, namely Parallel Rewiring and Turn Reduction. Parallel rewiring operates as a post-processing procedure to enhance an ESTC solution, while turn reduction introduces a secondary objective during the execution of Alg. 1.
Parallel Rewiring: As the original STC only explores the solution space of circumnavigating coverage paths, we develop two parallel rewiring procedures to refine any ESTC solution $\pi$ with non-circumnavigating local paths. While they do not theoretically improve the upper bound on the suboptimality of ESTC in Theorem 5, empirical results (see Sec. IV-A) often indicate an improvement in solution quality in practice. Type-A Parallel Rewiring searches for every subsequence in $\pi$ matching the pattern $(...,v,u,...,v,t,s,...,t,...)$ and collapses it into $(...,v,u,...,s,...t,...)$ if the resulting cost difference $w_{(u,s)}-[w_{(u,v)}+w_{(v,t)}+w_{(t,s)}]<0$. As shown in Fig. 5-(a), each collapsing operation effectively removes the intermediate vertices $v,t\in V$ without losing any vertex coverage. Type-B parallel rewiring identifies every pair of two vertically (or horizontally) parallel edges $(u_{1},u_{2}),(v_{1},v_{2})\in E(\pi)$ where $(u_{1},u_{3}),(u_{2},u_{4}),(s_{1},u_{3}),(s_{2},u_{4})\in E(\pi)$ are aligned horizontally (or vertically). Fig. 5-(b) shows that it removes the edges $(u_{1},u_{3}),(u_{2},u_{4}),(v_{1},v_{2})$ and inserts the edges $(u_{3},u_{4}),(u_{1},v_{1}),(u_{2},v_{2})$ if the cost difference after rewiring is less than 00, that is, $[w_{(u_{3},u_{4})}+w_{(u_{1},v_{1})}+w_{(u_{2},v_{2})}]-[w_{(u_{1},u_{3})}+w_{(u_{2},u_{4})}+w_{(v_{1},v_{2})}]<0$. Both types have a runtime complexity of $O(|V|^{2})$ with a naive implementation of two-layer iterations on $V$.
Turn Reduction: ESTC can be seamlessly integrated with a secondary objective of turn reduction in addition to its primary objective of path cost minimization. Specifically, we consider aligning the hyperedges when constructing the MST for ESTC. As the hypergraph $H=(V_{h},E_{h})$ is a four-way-connected 2d grid graph, we can categorize the hyperedges from $E_{h}$ by their directions as vertical and horizontal. Given a goal direction to align with, we prioritize the hyperedges traversal in Kruskal’s algorithm (line 1 of Alg. 1) by (1) their directions are identical to the goal direction and (2) the sum of their two endpoints’ degrees is smaller. The first prioritization is straightforward to align hyperedges with the goal direction as much as possible, whereas the second prioritization tries to first grow the MST into those less connected regions of $H$ for better alignment. As it only introduces two priority keys when constructing the MST, ESTC with turn reduction has the same time complexity as demonstrated in Theorem 6. Fig. 6 shows an example of applying turn reduction on an ESTC coverage path.
Figure 5: Local views of ESTC paths before and after parallel rewiring.
IV-D Integrating ESTC into Existing Grid-based MCPP Methods
While existing grid-based MCPP methods mostly focus on the special case of MCPP with strict assumptions of complete hypervertices and vertex-weighted graphs, we want to emphasize our more general formulated MCPP in Problem LABEL:problem:_mcpp is relatively unexplored in the MCPP literature, despite its significant practical applications. The proposed ESTC can be seamlessly integrated into these methods, allowing them effectively solve MCPP without the aforementioned assumptions. Considering MCPP for MRS $(G,I,R)$, we show several examples of adapting them for MCPP. Comprehensive numerical results for solving MCPP using these methods are provided in Sec. VII.
@@ IV-D Integrating ESTC into Existing Grid-based MCPP Methods
VOR: VOR is a naive baseline method introduced in [tang2024large]. It first decomposes $G$ into $k$ subgraphs $\{G_{i}\}_{i=1}^{k}$ as a Voronoi diagram [aurenhammer2000voronoi], with the set $R$ of root vertices as the pivot points and the shortest path distances on $G$ as the distance metric. Then, it generates coverage path $\pi_{i}$ on each subgraph $G_{i}$ via ESTC, constituting a solution for MCPP.
MFC: Multi-robot Forest Coverage (MFC)[zheng2007robot, zheng2010multirobot] heuristically computes a set $\{T_{i}\}_{i=1}^{k}$ of $k$ covering subtrees for the hypergraph $H$ using rooted-tree-cover[even2004min], where each hypervertex is included in at least one subtree. It then generates the coverage path on each $T_{i}$ via STC. To adapt it for MCPP, we make two modifications: (1) hypergraph $H$ is constructed as per Sec. IV-A instead of hypervertex-weighted in the original MFC and (2) STC is replaced with ESTC.
MIP: MIP [tang2023mixed] follows a similar two-level approach as MFC with one key difference. Instead of using rooted-tree-cover to find the subtree cover, it solves a mixed-integer program to find an optimal subtree cover that minimize the maximal subtree weight. By applying the same modifications to MIP as MFC, it can be used to solve MCPP.
MSTC: MSTC [tang2021mstc] is an improved version of MSTC [hazon2005redundancy], which operates differently from the above approach. MSTC first computes a single-robot coverage path via STC, then segments the path into $k$ segments each corresponding to a robot. MSTC then uses a shortest path algorithm to connect the root vertex with the cut points for each robot, forming the final coverage path for each robot. By replacing STC with ESTC, we can adapt MSTC to solve MCPP.
Figure 6: ESTC coverage path before (left) and after (right) turn reduction with the edges prioritized to align horizontally when constructing the MST.

### v2

@@ 2 $M\leftarrow$ find MST of $H$ via Kruskal’s algorithm
3 $\pi_{\delta}\leftarrow$ initial self-connecting local path for each $\delta\in V_{h}$
4 $\pi\leftarrow$ an empty path
5 for $\varepsilon=(\delta_{u},\delta_{v})\in$ the set of hyperedges in $M$ do
6       connect $\pi_{\delta_{u}},\pi_{\delta_{v}}$ via rerouting rule and add it to $\pi$
7
8return ESTC coverage path $\pi$
@@ 4 $\pi\leftarrow$ an empty path
Hyperedge Weights: The ESTC algorithm has a good property: Given any two spanning trees $T$ and $T^{\prime}$ with their resulting ESTC paths $\pi$ and $\pi^{\prime}$, respectively, their path difference (represented using the multisets of their edges) is a constant, given as follows:
$\begin{split}\mathcal{E}(\pi)\setminus\mathcal{E}(\pi^{\prime})=\bigcup_{\varepsilon\in\Delta E_{h}}E^{+}_{\varepsilon}\cup\bigcup_{\varepsilon^{\prime}\in\Delta E^{\prime}_{h}}E^{-}_{\varepsilon^{\prime}}&\quad\text{and}\\ \mathcal{E}(\pi^{\prime})\setminus\mathcal{E}(\pi)=\bigcup_{\varepsilon^{\prime}\in\Delta E^{\prime}_{h}}E^{+}_{\varepsilon^{\prime}}\cup\bigcup_{\varepsilon\in\Delta E_{h}}E^{-}_{\varepsilon}&,\end{split}$ (1)
where $\Delta E_{h}=E_{h}(T)\setminus E_{h}(T^{\prime})$ and $\Delta E^{\prime}_{h}=E_{h}(T^{\prime})\setminus E_{h}(T)$ represent the two hyperedge difference sets between $T$ and $T^{\prime}$. We then define the weight $w_{\varepsilon}$ for hyperedge $\varepsilon$ as below:
$w_{\varepsilon}=\sum_{e\in E^{+}_{\varepsilon}}w_{e}-\sum_{e\in E^{-}_{\varepsilon}}w_{e}.$ (2)
The definition of $w_{\varepsilon}$ implicitly encodes the ESTC path cost difference of whether $\varepsilon$ exists in the spanning tree. In Theorem 5, we validate that with the defined hyperedge weights in Eqn. (2), an MST of hypergraph $H$ always results in an optimal ESTC circumnavigating coverage path on $G$.
@@ IV-B Theoretical Analysis
ESTC is guaranteed to generate a path that covers all vertices of a connected grid graph $G$ based on the argument used in the proof of Lemma 3.1 in [55] since the argument is not affected by the addition of edge weights.
@@
ESTC achieves complete coverage on a connected grid graph $G$.
We now analyze the time complexity of ESTC below.
@@
The time complexity of ESTC (Alg. 1) on a grid graph $G=(V,E)$ is $O(|V|\log|V|)$.
Proof.
To construct the hypergraph $H=(V_{h},E_{h})$ from $G=(V,E)$ as described in Sec. IV-A, ESTC first iterates through $V$ to build $V_{h}$ and then through $V_{h}$ to build $E_{h}$. Given that $|V_{h}|$ is proportional to $|V|$, the time complexity of this construction is $O(|V|)$. To find the MST of $H$, ESTC uses Kruskal’s algorithm, which incurs a time complexity of $O(|E|\log|E|)$. Initializing the local paths and connecting them via the rerouting rule entail a time complexity of $O(|V|+|E|)$. As $O(|E|)=O(|V|)$ for grid graphs, the overall time complexity of ESTC combines to $O(|V|\log|V|)$. ∎
The following theorem validates that our definition of hyperedge weights ensures that the ESTC circumnavigating path derived from an MST of $H$ has the smallest cost among those derived from any spanning trees of $H$.
@@
The cost of the circumnavigating path derived from a minimum spanning tree (MST) of the hypergraph $H$ is no greater than that from any other spanning tree of $H$.
Proof.
For hypergraph $H=(V_{h},E_{h})$, consider its two spanning trees $T$ and $T^{\prime}$ and their corresponding ESTC circumnavigating paths $\pi$ and $\pi^{\prime}$, respectively. With Eqn. (1), the path cost difference $\Delta c$ between $\pi$ and $\pi^{\prime}$ can be computed as:
$\Delta c$ $=c(\pi)-c(\pi^{\prime})=\sum_{e\in\mathcal{E}(\pi)\setminus\mathcal{E}(\pi^{\prime})}w_{e}-\sum_{e\in\mathcal{E}(\pi^{\prime})\setminus\mathcal{E}(\pi)}w_{e}$
$=\sum_{\varepsilon\in\Delta E_{h}}\sum_{e\in E^{+}_{\varepsilon}}w_{e}+\sum_{\varepsilon^{\prime}\in\Delta E^{\prime}_{h}}\sum_{e\in E^{-}_{\varepsilon^{\prime}}}w_{e}$
$-\sum_{\varepsilon^{\prime}\in\Delta E^{\prime}_{h}}\sum_{e\in E^{+}_{\varepsilon^{\prime}}}w_{e}-\sum_{\varepsilon\in\Delta E_{h}}\sum_{e\in E^{-}_{\varepsilon}}w_{e}.$
With Eqn. (2) substituted into the above, we have:
$\begin{split}\Delta c&=\sum_{\varepsilon\in\Delta E_{h}}\left[\sum_{e\in E^{+}_{\varepsilon}}w_{e}-\sum_{e\in E^{-}_{\varepsilon}}w_{e}\right]\\ &-\sum_{\varepsilon^{\prime}\in\Delta E^{\prime}_{h}}\left[\sum_{e\in E^{+}_{\varepsilon^{\prime}}}w_{e}-\sum_{e\in E^{-}_{\varepsilon^{\prime}}}w_{e}\right]\\ &=\sum_{\varepsilon\in\Delta E_{h}}w_{\varepsilon}-\sum_{\varepsilon^{\prime}\in\Delta E^{\prime}_{h}}w_{\varepsilon^{\prime}}.\end{split}$ (3)
If $T$ is an MST and $T^{\prime}$ is an arbitrary spanning tree, we know that $\forall\varepsilon\in\Delta E_{h}\subseteq E_{h}(T)$ and $\varepsilon^{\prime}\in\Delta E^{\prime}_{h}\subseteq E_{h}(T^{\prime})$, we have $w_{\varepsilon}\leq w_{\varepsilon^{\prime}}$, where $E_{h}(T)$ and $E_{h}(T^{\prime})$ are the edge sets of $T$ and $T^{\prime}$, respectively. Substituting it into Eqn. (3), we have $\Delta c\leq 0$ for an MST $T$ and an arbitrary spanning tree $T^{\prime}$. ∎
The following theorem shows that the optimal circumnavigating path on $H$ yields an optimal CPP solution when all hypervertices are complete and a bounded suboptimal CPP solution in general.
Theorem 6.
ESTC returns a coverage path with a cost at most $2\cdot\frac{w_{\text{max}}}{w_{\text{min}}}\cdot(1+\frac{n_{c}-1}{|V|})$ times the cost of the optimal single-robot coverage path, where $n_{c}$ is the number of complete hypervertices and $w_{\text{max}}$ and $w_{\text{min}}$ are the maximum and minimum hyperedge weights, respectively.
Proof.
@@
Consider a CPP instance $(G,I,R)$, where $G$ is contracted into hypergraph $H=(V_{h},E_{h})$ with $n_{c}$ complete hypervertices. For any optimal valid coverage path $\pi^{*}$, it follows that $V(\pi)=V$ and both the first and last vertices of $\pi$ are the root vertex, implying $|\mathcal{E}(\pi^{*})|\geq|V|$, with equality only if there are no incomplete hypervertices in $H$. The suboptimality ratio $\rho$ for any ESTC path $\pi$ resulted from MST $T$ is given by:
$\rho=\frac{c(\pi)}{c(\pi^{*})}=\frac{\sum_{e\in\mathcal{E}(\pi)}w_{e}}{\sum_{e\in\mathcal{E}(\pi^{*})}w_{e}}\leq\frac{|\mathcal{E}(\pi)|\cdot w_{\text{max}}}{|\mathcal{E}(\pi^{*})|\cdot w_{\text{min}}}\leq\frac{|\mathcal{E}(\pi)|\cdot w_{\text{max}}}{|V|\cdot w_{\text{min}}}.$
We now categorize $V_{h}$ into four disjoint sets such that $V_{h}=\bigcup_{i=1}^{4}V_{h}^{(i)}$ and each $V_{h}^{(i)}$ contains hypervertices with $i$ vertices located in it. As illustrated in Fig. (4a), for any $\delta\in V_{h}^{(i)}$, the initial ESTC self-connecting local path $|\pi_{\delta}|=4$ if $i=4$ or $3$, $|\pi_{\delta}|=2$ if $i=2$, and $|\pi_{\delta}|=0$ if $i=1$. We categorize the hyperedge set $E_{h}(T)$ of MST $T$ into $E_{h}^{(2)}(T)$ and $E_{h}^{(1)}(T)$ according to the number of crossing edges as described in Fig. (4c) and (4d). By the initialization of the self-connecting local paths and the rerouting rule, the cardinality of the multiset $\mathcal{E}(\pi)$ can be computed as:
$|\mathcal{E}(\pi)|=4|V_{h}^{(4)}|+4|V_{h}^{(3)}|+2|V_{h}^{(2)}|+2|E_{h}^{(1)}(T)|.$ (4)
As $|V|=\sum_{i=1}^{4}i\cdot|V_{h}^{(i)}|$, the bound can be further refined as:
$\begin{split}\rho&\leq\frac{w_{\text{max}}}{w_{\text{min}}}\cdot\left[1+\frac{|V_{h}^{(3)}|-|V_{h}^{(1)}|+2|E_{h}^{(1)}(T)|}{|V|}\right]\\ \end{split}.$ (5)
Substituting $|E_{h}^{(1)}(T)|\leq|E_{h}(T)|=|V_{h}|-1=\sum_{i=1}^{4}|V_{h}^{(i)}|-1$ and $n_{c}=|V_{h}^{(4)}|$ into Eqn. (5), we obtain a simpler yet looser bound:
$\rho\leq 2\cdot\frac{w_{\text{max}}}{w_{\text{min}}}\cdot\left(1+\frac{n_{c}-1}{|V|}\right).$ (6)
@@
With the tighter bound in Eqn. (5), we can see that ESTC is optimal ($\rho=1$) when $G$ is unweighted ($\frac{w_{\text{max}}}{w_{\text{min}}}=1$) and all hypervertices are complete ($|V_{h}^{(3)}|=|V_{h}^{(1)}|=|E_{h}^{(1)}(T)|=0$). The bound in Eqn. (6) may appear relatively loose. However, the last multiplier $(1+\frac{n_{c}-1}{|V|})\in[1,2)$ as $n_{c}\leq|V|$. In fact, it can be seen that as $n_{c}$ approaches 00, the bound in Eqn. (6) gets tighter and closer to $2\cdot\frac{w_{\text{max}}}{w_{\text{min}}}$. This is because when $H$ has more incomplete hypervertices, the optimal solution $\pi^{*}$ necessitates a rerouting rule, such as that in Full-STC, to revisit edges through incomplete hypervertices to achieve complete coverage.
IV-C Local Optimizations
We introduce two efficient local optimizations for ESTC, namely Parallel Rewiring and Turn Reduction. Parallel Rewiring operates as a post-processing procedure to refine the coverage path produced by ESTC, by strategically adjusting path segments locally to reduce the path cost. Turn Reduction serves as a secondary objective during the execution of ESTC (see Alg. 1) by reducing the number of turns in the path, which is particularly beneficial in environments where turning is costly or undesirable. While these local optimizations do not theoretically improve the upper bound on the suboptimality of ESTC in Theorem 6, empirical observations (see Sec. IV-A) often demonstrate a notable improvement in solution quality.
Parallel Rewiring: As the original STC only explores the solution space of circumnavigating coverage paths, we develop two parallel rewiring procedures to refine any ESTC solution $\pi$ using non-circumnavigating local path segments. Type-A Parallel Rewiring searches for every subsequence in $\pi$ matching the pattern $(\ldots,v,u,\ldots,v,t,s,\ldots,t,\ldots)$ and collapses it into $(\ldots,v,u,\ldots,s,\ldots t,\ldots)$ if the resulting cost difference $w_{(u,s)}-[w_{(u,v)}+w_{(v,t)}+w_{(t,s)}]<0$. As shown in Fig. (6a), each collapsing operation effectively removes the intermediate vertices $v,t\in V$ without losing any vertex coverage. Type-B parallel rewiring identifies every pair of two vertically (or horizontally) parallel edges $(u_{1},u_{2}),(v_{1},v_{2})\in\mathcal{E}(\pi)$ where $(u_{1},u_{3}),(u_{2},u_{4}),(s_{1},u_{3}),(s_{2},u_{4})\in\mathcal{E}(\pi)$ are aligned horizontally (or vertically). As shown in Fig. (6b), it then replaces the edges $(u_{1},u_{3}),(u_{2},u_{4}),(v_{1},v_{2})$ with $(u_{3},u_{4}),(u_{1},v_{1}),(u_{2},v_{2})$ if the resulting cost difference $[w_{(u_{3},u_{4})}+w_{(u_{1},v_{1})}+w_{(u_{2},v_{2})}]-[w_{(u_{1},u_{3})}+w_{(u_{2},u_{4})}+w_{(v_{1},v_{2})}]<0$. Both types have a time complexity of $O(|V|^{2})$ with a naive implementation involving two-layer nested iterations through $V$.
Turn Reduction: ESTC can be seamlessly integrated with a secondary objective of turn reduction in addition to its primary objective of path cost minimization. Specifically, when ESTC constructing the MST, it prioritizes the alignment of hyperedges. As the hypergraph $H=(V_{h},E_{h})$ is a four-neighbor 2d grid, its hyperedges can be categorized by their orientation as vertical or horizontal. To align with a given target orientation, the ESTC prioritizes the hyperedges traversal in Kruskal’s algorithm (line 1 of Alg. 1) to favor a hyperedge where (1) its orientation matches the given target orientation and (2) sum of the degrees of its two endpoints is the smallest. The first prioritization ensures orientation consistency as much as possible, whereas the second prioritization aims to extend the MST into less connected regions of $H$ for better alignment. Turn Reduction requires adding only two priority keys when constructing the MST and thus maintains the same time complexity for ESTC as demonstrated in Theorem 4. Fig. (6c) shows an example of applying Turn Reduction to an ESTC coverage path.
Figure 6: ESTC coverage path before (top row) and after local optimizations (bottom row). (a)(b) The two types of parallel rewiring. (c) Turn reduction with horizontally aligned hyperedges.
IV-D Integration into Existing MCPP Methods
Existing grid-based MCPP methods often focus on special cases of MCPP that assume complete hypervertices and vertex-weighted graphs. However, our more generally formulated MCPP, as defined in Problem 1, addresses a broader scope that remains relatively unexplored in the literature, despite its significant practical relevance. The proposed ESTC paradigm can be seamlessly integrated into these existing methods, allowing them to effectively solve MCPP even without the above constraints typically assumed. To demonstrate how existing MCPP methods can be adapted using the ESTC paradigm to solve an MCPP instance on any grid graph $G$, even when some hypervertices are incomplete, we use the following examples categorized into multi-tree and single-tree methods.
Figure 7: The flowchart of our LS-MCPP algorithmic framework for instance floor-small. Given an initial MCPP solution $\Pi_{0}$, LS-MCPP iteratively samples a pool $O$ from three pools of dedicated operators and then samples an operator $o$ from $O$, to transform a set $\mathcal{G}$ of subgraphs from one to another. It then applies ESTC on the subgraphs to evaluate the new MCPP solution $\Pi$, which is adaptively accepted according to its makespan $\theta$. A periodic subroutine of forced deduplication is called at the end of each iteration, and the best solution $\Pi^{*}$ is recorded if $\Pi$ is better with a lower makespan.
IV-D1 Multi-Tree Methods
For MCPP instance $(G,I,R)$, we first generate a set $\{G_{i}=(V_{i},E_{i})\}_{i=1}^{k}$ of $k$ connected subgraphs of $G$. Each $G_{i}$ contains the root vertex $r_{i}$. We then solve each sub-CPP instance for robot $i$ on $G_{i}$ as described in Alg. 1 to generate its coverage path. The following theorem shows a sufficient condition for complete coverage.
@@ IV-D1 Multi-Tree Methods
If each $G_{i}$ is connected and $\bigcup_{i=1}^{k}V_{i}=V$, then using ESTC is guaranteed to achieve complete coverage of $G$.
Proof.
Given that $G_{i}$ is connected, ESTC is guaranteed to generate a path that covers all its vertices due to Theorem 3. If $\bigcup_{i=1}^{k}V_{i}=V$, then the coverage paths generated for all $G_{i}$ jointly cover all vertices of $G$. ∎
The design of our LS-MCPP framework is based on the principle expounded in the above theorem. It aims to search for a good set of subgraphs $G_{i}$ for the robots, each being connected and containing the root vertex $r_{i}$, with the property that the union of all their vertices comprises the entire vertex set of $G$ for complete coverage. It now suffices to outline the details below for adapting existing methods to compute the $k$ subgraphs, with comprehensive numerical results in Sec. VII.
@@
VOR: VOR is a baseline method introduced in [21] that is inherently compatible with ESTC. It partitions $G$ into $k$ subgraphs $\{G_{i}\}_{i=1}^{k}$ as a Voronoi diagram [57], using the set $R$ of root vertices as the pivot points and shortest path distances on $G$ as the distance metric.
MFC: Multi-Robot Forest Coverage (MFC) [12, 13] employs the Rooted-Tree-Cover procedure [26] to heuristically compute a rooted tree cover on the quadrant coarsened grid $\mathcal{H}$—a forest of $k$ trees $\{T_{i}\}_{i=1}^{k}$, each rooted at the hypervertex containing the root vertex $r_{i}$, which jointly cover all hypervertices of $\mathcal{H}$. MFC then generates a coverage path for each robot $i$ by applying STC on its tree $T_{i}$. To integrate MFC with ESTC, we apply the Rooted-Tree-Cover procedure on the hypergraph $H$ (constructed as per Sec. IV-A) instead of $\mathcal{H}$ to generate a rooted tree cover on $H$ that induces the set of $k$ subgraphs $\{G_{i}\}_{i=1}^{k}$ of $G$.
MIP: MIP [15] solves a mixed-integer program to compute an optimal rooted tree cover on the quadrant coarsened grid $\mathcal{H}$ that minimizes the maximum tree weight. To integrate MIP with ESTC, we adapt the mixed-integer program to compute an optimal rooted tree cover on the hypergraph $H$ (constructed as per Sec. IV-A) instead of $\mathcal{H}$, which induces the set of $k$ subgraphs $\{G_{i}\}_{i=1}^{k}$ of $G$.
IV-D2 Single-Tree Methods
Single-tree methods, unlike multi-tree methods, apply the STC paradigm to only a single tree: STC first computes an MST on the quadrant coarsened grid $\mathcal{H}$ and subsequently generates a coverage path that circumnavigates the MST. The path is then partitioned into $k$ segments, one for each robot. MSTC [11] partitions the path such that each segment starts at a unique root vertex. Its improvement MSTC [14] greedily identifies balanced partitioning points along the path to potentially reduce the makespan and then connects each root vertex to its respective segment via shortest paths. We can adapt MSTC to solve MCPP for scenarios with incomplete hypervertices by simply replacing STC with ESTC.